OpenStudy (anonymous):

solve the equation. check your answer. 4y+5=-31

OpenStudy (anonymous):

4y+5=-31 4y=-31-5 4y=-36 y=-36/4 =-9

OpenStudy (anonymous):

i solved the equation i just need to know how to check it or plug it back in

OpenStudy (jannine24):

take your -9 then plug it back in for 4

OpenStudy (anonymous):

Plug -9 in for y.

OpenStudy (jannine24):



OpenStudy (anonymous):

so, does 4(-9)+5 equal -31?

OpenStudy (jannine24):

yes

OpenStudy (anonymous):

im a bit confused but i think i got it?

OpenStudy (jannine24):

what r u confused on

OpenStudy (anonymous):

just a bit on how you worked it, im trying to understand that part

OpenStudy (jannine24):

u r solving for the variable y

OpenStudy (anonymous):

get common terms together..

OpenStudy (anonymous):

thank you all the help!